# Robustly overfitting latents for flexible neural image compression

### v1

@@ 4.3 Experimental results
Table 2 represents the interpolation between different functions, expressed in true R-D loss. Values for $a<1$ indicate methods that tend to have (extremely) large gradients for $v$ close to the corners, while high values of $a$ represent a method that tends to a (reversed) step function. We find that SLL with small $a=\{0.01,0.3\}$ diverges and results in large loss values compared to the rest. Additionally, the loss curves show unstable behavior for these functions, which can be found in Appendix A.1. For $a=\{0.8,\cdots,2.25\}$, the loss shows stable behavior in terms of its curves and final loss at $t=2000$. Furthermore, for $a=1.65$ (approximately $\operatorname{atanh}$), the method shows non-optimal behavior due to a slightly unstable loss curve, see Appendix A.1, this may be due to setting $\tau_{max}=1$ instead of optimal setting $\tau_{max}=[0.4,0.5]$. Note that the difference in loss $0.7410$ for SSL $a=0.65$ and $0.7418$ for $\operatorname{atanh}$ with $\tau_{max}=1$ (see Table 1). This difference may be because SSL is an approximation to the $\operatorname{atanh}$ function and not exactly equal.
Three-class rounding
In Table 3, the true R-D loss for two versus three-class rounding can be found at iteration $t=500$ and in brackets $t=2000$ iterations. For each method, we performed a grid search over the hyperparameters $r$ and $n$. Additionally, for the extended SSL, we also performed a grid search over $a$ and found the best setting to be $a=1.4$. As can be seen in the table, most impact is made with the extended version of the linear of SGA+, in terms of the difference between the two versus three-class rounding at iteration $t=500$ with loss difference $0.0035$ and $t=2000$ with $0.0006$ difference. There is a small difference at $t=500$ for the extended cosine version. In general, we find that running models longer results in convergence to similar values. SSL converges to equal values for two- and three-class rounding. This makes three-class rounding attractive under a constraint optimization budget, possibly because it is easier to jump between classes.
Table 3: True R-D loss of two versus three-class rounding for SGA+ with the extended version of the linear, cosine, and SSL method at iteration 500 and in brackets after 2000 iterations.
@@
4.4 Tecnick
To test how our method performs on another dataset, we use the Tecnick dataset (Asuni & Giachetti, 2014). We run baselines $\operatorname{atanh}$ and the base model and compare against SSL with $a=\frac{4}{3}$. Figure 3 shows the corresponding R-D curves, using image quality metric PSNR versus BPP, after $t=500$ iterations. We find that both refining methods greatly improve the compression performance in terms of the R-D trade-off. Additionally, our proposed method outperforms the baselines and shows how it boosts performance, especially for the smallest and highest $\lambda=\{0.001,0.08\}$. The R-D plot after $t=2000$ iterations can be found in Appendix A.2.
Figure 3: Rate-distortion performance on the Tecnick dataset at iteration 500 for the base model, $\operatorname{atanh}$ and SSL.
4.5 Semi-multi-rate behavior
Figure 4: RD-plot with semi-multi-rate behavior.
An interesting observation is that one does not need to use the same $\lambda$ during refinement of the latents, as used during training. As a consequence of this approach, we can optimize to a neighborhood of the R-D curve without the need to train a new model from scratch. For instance, we only need three out of seven models to have better performance over the entire curve compared to the base model. We experimented with different values for $\lambda$. For every pre-trained model, we ran our best SSL approach using $\lambda\in\{0.0001,0.0005,0.001,0.0025,0.005,0.01,0.02,0.04,\newline 0.08,0.16,0.32\}$ to obtain a semi-multi-rate curve. In Figure 4, we have plotted the R-D curve of the base model (lime green line) and its corresponding R-D curves, obtained when refining the latents with the proposed $\lambda$’s. We pruned away points that fell too far below the base curve and refer to Appendix A.3 for the unpruned version.
(a) Original Kodak Image
(b) Base Model $\lambda=0.001$
(0.08 BPP vs 27.08 PSNR)
(c) Base model $\lambda=0.0025$, SSL trained with $\lambda=0.0005$
(0.08 BPP vs 27.73 PSNR)
Figure 5: Qualitative comparison of an Kodak image which can be best viewed electronically. Base model for $\lambda=0.001$ compared to base model $\lambda=0.0025$ tuned with $\lambda=0.0005$ during the refinement procedure for SSL.
Qualitative results
In Figure 5, we demonstrate the effects of using the semi-multi-rate strategy. We compare the original image, the compressed image using the base model for $\lambda=0.001$, and the compressed image using SSL with a base $\lambda=0.0025$ and a target $\lambda=0.0005$. We have chosen the latter two to compare as their BPP values are similar, while the base models are different. For the image compressed by SSL, we observe that there are less artifacts visible. For instance, looking at the fence we see more texture compared to the base model. Additionally, looking at the sky we find less artifacts compared to the base model.
4.6 Hyperparameter settings
Similar to (Yang et al., 2020), we find an optimal learning rate of $0.005$ for $\operatorname{atanh}$ and find the best performance for STE with a smaller learning rate of $0.0001$, yet STE still has trouble converging. Additionally, we find the best performance for $\operatorname{atanh}$ with $\tau_{max}\in[0.4,0.5]$. For SGA+, we use optimal convergence settings, which are a fixed learning rate of $0.005$, a temperature rate of $c=0.001$, and $\tau_{max}=1$. Furthermore, experimentally, we find approximately equal and best performance for SLL with $a\in[1.3,1.4]$. For SSL we choose $a=\frac{4}{3}$, see corresponding probability in Figure 0(a).
5 Discussion and conclusion
Training neural image compression models is a time-consuming and difficult task. In practice, finding optimal encoding comes with difficulties and can lead to sub-optimal performance of these networks. Refining latents of a pre-trained network has been shown to improve the compression performance by adding uniform noise or using a stochastic method called SGA (Campos et al., 2019; Yang et al., 2020).
We extend this idea and propose SGA+. We show how SGA+ has nicer properties, which aids the compression performance. We introduced SSL that can approximately interpolate between all of the proposed methods. Further, we show how our best-performing method SSL outperforms the baselines and that it is more stable under varying conditions. We give a general notation and demonstrate how the extension to three-class rounding improves the convergence of the SGA+ methods but comes with the cost of fine-tuning extra hyperparameters. Finally, we show how our proposed method performs when refining latents on the Tecnick dataset and how we can use refining of the latents to obtain semi-multi-rate behavior.
In general, it applies for each method that as the temperature rate has reached a stable setting, the longer you train, the closer together the performance will be. However, when under a constrained optimization budget, running SGA+ shorter already gives good results. The best results are obtained while tuning the hyperparameter of SSL. Further, our experiments show that the linear version for SGA+ is least sensitive to hyperparameter changes.
Broader impact
The development and improvement of neural compression techniques play an important role in our increasingly data-driven society. Being able to learn compression codecs as opposed to developing them saves time and effort. There is a rich array of signal-based data types available that benefit from compression. A smaller data footprint reduces the requirements for storage, networking, and computing resources. It thus also means smaller investments in terms of raw resources, manufacturing, energy, and labor.
Neural image compression methods still require a large amount of computation to be trained. This process has tremendous adverse effects on the environment, as established earlier. Our work aims to make these models more flexible through the proposal of methods that improve on these already trained models. Improved flexibility can result in improved reusability, which may reduce the environmental impact of neural image codecs.
Appendix A Additional results
In this appendix, additional experiment results can be found.
A.1 Interpolation
Table 2 presents the true R-D loss results for the interpolation with different $a$ settings for SSL. In Figure 6, the corresponding overall performance of the methods can be found. As can be seen in Figure 5(a), for $a=\{0.01,0.30\}$, the functions diverge, resulting in large loss values. For $a=0.65$, we find that the loss curve is slightly unstable at the beginning of training, which can be seen in the bending of the curve, indicating non-optimal settings. This may be due to the fact that we run all methods with the same $\tau_{max}=1$ for a fair comparison. Additionally, note that SSL with $a=0.65$ obtains a true R-D loss of $0.7410$ compared to $0.7418$ for $\operatorname{atanh}$ with the same settings. This is due to the fact that SSL, especially in the tails of the probability, is slightly more straight-curved compared to the $\operatorname{atanh}$ when looking at its probability space.
@@ A.1 Interpolation
Figure 6: Interpolation performance plots of different $a$ settings for SLL (a) True R-D Loss (b) Difference in loss (c) PSNR (d) BPP
A.2 Rate-Distortion performance
We evaluate our best performing method SSL with $a=\frac{4}{3}$ on the Kodak and Tecnick datasets, by computing the R-D performance, average over each of the datasets. The R-D curves use image quality metric PSNR versus BPP on the Kodak and Tecnick dataset. Recall that as base model we use the pre-trained mean-scale hyperprior, trained with $\lambda=\{0.001,0.0025,0.005,$ 0.01,0.02,0.04,0.08}0.01,0.02,0.04,0.08\}0.01 , 0.02 , 0.04 , 0.08 }.
Kodak
Figure 7 shows the R-D curve for refining the latents, evaluated on Kodak. We compare SLL against baselines: STE, uniform noise, $\operatorname{atanh}$ and the base model at iteration $t=500$ (see Figure 6(a)) and after full convergence at $t=2000$ (see Figure 6(b)). As can be seen in Figure 6(a), STE performs slightly better than the base model, while after $t=2000$ iterations the method performs worse, this also reflected in the corresponding true loss curve for $\lambda=0.01$ (see Figure 1(a)), which diverges. Remarkably, for the smallest $\lambda=0.001$, STE performs better than at $t=500$. Adding uniform noise results in better performance when running the method longer. Further, SSL outperforms $\operatorname{atanh}$ slightly, while the difference is more pronounced when running for $t=500$ iterations.
(a) R-D curve at iteration $t=500$
(b) R-D curve at iteration $t=2000$
Figure 7: R-D performance on Kodak of the base model, STE, Uniform noise, SGA $\operatorname{atanh}$ and SSL with $a=\frac{4}{3}$
Tecnick
Figure 8 shows the R-D curve when refining latents on the Tecnick dataset, after $t=2000$ iterations. As can be seen in the plot, we find that the longer the methods run, the closer the performance lies to each other. When there is a limited budget available, one can run the refinement process for $t=500$ iterations.
Figure 8: R-D performance on Tecnick after $t=2000$ iterations, which compares SSL with the baselines: base model and $\operatorname{atanh}$.
A.3 Semi multi-rate behavior
Figure 9: R-D performance with semi-multi-rate behavior. For each base model, we ran SGA+ with $\lambda\in\{0.0001,0.0005,\dots,0.16,0.32\}$ and depicted the resulting BPP/PSNR point.
Figure 9 shows the different R-D curves when refining the latents using different values for $\lambda$. For each model trained using $\lambda\in\{0.001,0.0025,0.005,0.01,0.02,0.04,0.08\}$, we run SSL with $a=\frac{4}{3}$ for $t=2000$ iterations for all $\lambda\in\{0.0001,0.0005,0.001,0.0025,0.005,0.01,0.02,0.04,0.08,0.16,0.32\}$. We depicted the base curve alongside the curves for each base model. We observe that using SGA+ in this manner gives samples diverse enough such that there is always a point on the resulting curve that is better than the base curve. As a result, latent fine-tuning allows for semi-multi-rate behavior. We also observed that when the values for $\lambda$ used between training and SGA+ differ too much, the resulting BPP versus PSNR point becomes worse than the base model.

### v3

@@ Appendix C Mean-Scale Hyperprior
Implementations details
The pre-trained mean-scale hyperpriors are trained from scratch on the full-size CLIC 2020 Mobile dataset Toderici et al. (2020), mixed with the ImageNet 2012 dataset Russakovsky et al. (2015) with randomly cropped image patches taken of size $256\times 256$. For ImageNet, only images with a size larger than $256$ for height and width are used to prevent bilinear up-sampling that negatively affects the model performance. During training, each model is evaluated on the Kodak dataset Kodak . The models were trained with $\lambda=\{0.001,0.0025,0.005,$ 0.01,0.02,0.04,0.08}0.01,0.02,0.04,0.08\}0.01 , 0.02 , 0.04 , 0.08 }, with a batch size of 32 and Adam optimizer with a learning rate set to $1e^{-4}$. The models are trained for 2M steps, except for model $\lambda=0.001$, which is trained for 1M steps and model $\lambda=0.08$, which is trained for 3M steps. Training runs took half a week for the 1M step model, around a week for the 2M step models, and around 1.5 weeks for the larger 3M step model. We ran all models and methods on a single NVIDIA A100 GPU. Further, the models for $\lambda=\{0.04,0.08\}$ are trained with $256$ hidden channels and the model for $\lambda=0.001$ is trained with $128$ hidden channels. The remaining models are trained with hidden channels set to $192$.
C.1 R-D Performance
We evaluate our best-performing method SSL on the Kodak and Tecnick datasets, by computing the R-D performance, average over each of the datasets. The R-D curves use image quality metric PSNR versus BPP on the Kodak and Tecnick dataset. Recall that as base model we use the pre-trained mean-scale hyperprior, trained with $\lambda=\{0.001,0.0025,0.005,$ 0.01,0.02,0.04,0.08}0.01,0.02,0.04,0.08\}0.01 , 0.02 , 0.04 , 0.08 }. For SSL we choose $a=\frac{4}{3}$ as we found that this setting achieves the best R-D loss overall at 2000 iterations. This is a lower setting compared to the model by Cheng et al. (2020). The hyperparameters are similar to what we reported for Cheng et al. (2020) but with two main differences. We found that we could increase the learning rate by a factor $10$ to $0.005$ for $\operatorname{atanh}$ and SGA+. We also found that a lower $a\in[1.3,1.4]$ was optimal.
Kodak
Figure C.2 shows the R-D curve for refining the latents, evaluated on Kodak. We compare SLL against baselines: STE, uniform noise, $\operatorname{atanh}$ and the base model at iteration $t=500$ (see Figure 2(a)) and after full convergence at $t=2000$ (see Figure 2(b)). As can be seen in Figure 2(a), STE performs slightly better than the base model, while after $t=2000$ iterations the method performs worse, this is also reflected in the corresponding true loss curve for $\lambda=0.01$ (see Figure 1(a)), which diverges. Remarkably, for the smallest $\lambda=0.001$, STE performs better than at $t=500$. Adding uniform noise results in better performance when running the method longer. Comparing the R-D curves, Figure B.1 to Figure C.2, we find that most impact is made at $t=500$ iterations. However, for the model similar to Yang et al. (2020) the performance lay closer to the performance of $\operatorname{atanh}$, than for the model of Cheng et al. (2020).
(a) True R-D loss
(b) Difference loss
(c) PSNR
(d) BPP
Figure C.1: Performance plots of (a) True R-D Loss (b) Difference in loss (c) PSNR (d) BPP.
(a) R-D curve at iteration $t=500$
(b) R-D curve at iteration $t=2000$
Figure C.2: R-D performance on Kodak of the base mean-scale hyperprior model, STE, Uniform noise, SGA $\operatorname{atanh}$ and SSL with $a=\frac{4}{3}$.
Tecnick
Figure C.3 shows the R-D curve when refining latents on the Tecnick dataset, after $t=500$ and $t=2000$ iterations. As can be seen in the plot, we find that the longer the methods run, the closer the performance lies to each other. The improvement by SSL compared to $\operatorname{atanh}$ is greater for Tecnick than for Kodak.
(a) R-D curve at iteration $t=500$
(b) R-D curve at iteration $t=2000$
Figure C.3: R-D performance on Tecnick of the base mean-scale hyperprior model, SGA $\operatorname{atanh}$ and SSL with $a=\frac{4}{3}$. Best viewed electronically.
CLIC
(a) R-D curve at iteration $t=500$
(b) R-D curve at iteration $t=2000$
Figure C.4: R-D performance on CLIC of the base mean-scale hyperprior model, SGA $\operatorname{atanh}$ and SSL with $a=\frac{4}{3}$. Best viewed electronically.
(a) R-D semi-multi-rate curve at iteration $t=500$
(b) R-D semi-multi-rate curve at iteration $t=2000$
Figure C.5: R-D performance on Kodak of the base mean-scale hyperprior model, SGA $\operatorname{atanh}$ and SSL with $a=\frac{4}{3}$. Each point is optimized with a different target $\lambda\in\{0.001,0.0025,0.005,0.01,0.02,0.04,0.08\}$.
Figure C.4 shows the R-D curve when refining latents on the CLIC dataset, after $t=500$ and $t=2000$ iterations. Similar to the previous results, we find that the longer the methods run, the closer the performance lies to each other and that running the method shorter already gives better performance, compared to the base model.
Table C.2: Pairwise Comparison of BD-PSNR and BD-Rate for the Kodak, Tecnick, and CLIC dataset on the Mean-Scale Hyperprior Model.
BD-PSNR (dB) BD-Rate (%)
500 steps 2000 steps 500 steps 2000 steps
Kodak Tecnick CLIC Kodak Tecnick CLIC Kodak Tecnick CLIC Kodak Tecnick CLIC
Base vs SSL 0.68 1.21 1.03 0.91 1.50 1.33 -13.52 -22.77 -21.87 -17.69 -28.17 -27.86
Base vs Atanh 0.59 1.06 0.89 0.87 1.46 1.30 -11.90 -20.20 -19.14 -17.03 -27.49 -27.28
Atanh vs SSL 0.09 0.15 0.14 0.04 0.04 0.03 -1.80 -3.22 -3.14 -0.82 -1.03 -0.74
BD-Rate Gain
In Table C.2, we computed the change in BD-PSNR and BD-rate for the mean-scale hyperprior model. We observe that SSL is slightly better than atanh, although the difference between them is smaller than reported on the Cheng et al. (2020) model. The gap between 500 steps and 2000 steps is also smaller compared to the results in 1.
C.2 Analysis
In this appendix, we analyze additional experiments for the model, similar to those in Yang et al. (2020). The results for the analysis are obtained from a pre-trained model trained with $\lambda=0.01$.
Table C.3: True R-D loss for different learning settings of: $\operatorname{atanh}$ and SSL with $a=\frac{4}{3}$. At $t=2000$ iterations and in brackets $t=500$ iterations
[…]
0.02 $0.7386$ ($0.7506$) $0.7491$ ($0.7627$)
0.01 $0.7375$ ($0.7498$) $0.7411$ ($0.7540$)
0.005 (base) $0.7380$ ($0.7521$) $0.7408$ ($0.7570$)
Learning rates
We run SSL and $\operatorname{atanh}$ with higher learning rate settings of $0.02$ and $0.01$ and compare it to the results obtained with learning rate $0.005$. Figure C.6 shows the corresponding loss curves and Table C.3 shows the corresponding loss values at $t=\{500,2000\}$ iterations. We find that for a learning rate of $0.01$ the gap between $\operatorname{atanh}$ versus SSL at 500 iterations is only around $14.3\%$ smaller and it remains pronounced, while with a learning rate of $0.01$ the gap at 2000 iterations is around $28.6\%$ better. This concludes that SSL benefits more than $\operatorname{atanh}$ at 2000 iterations, with a higher learning rate. More interestingly, for a learning rate of $0.02$, $\operatorname{atanh}$ diverges whereas SSL still reaches comparable performance. This highlights the sensitivity of $\operatorname{atanh}$.
Figure C.6: True R-D loss curves for different learning rates settings for method SSL and $\operatorname{atanh}$.
Table C.4: True R-D loss for different $\tau_{max}$ settings of: $\operatorname{atanh}(v)$, linear, cosine and SSL with $a=\frac{4}{3}$. Lowest R-D loss per column is marked with: $\downarrow$. Note that the function containing $\operatorname{atanh}$ is unnormalized.
Function \$\tau_{max}$ 0.2 0.4 0.6 0.8 1.0
$\cos^{2}(\frac{v\pi}{2})$ $0.7496$ $0.7414$ $0.7393$ $0.7387$ $0.7384$
$\exp{\operatorname{atanh}(v)}$ 00 $0.0002$ $0.0022$ $0.0033$ $0.0038$
$1-v$ (linear) $0130.0$ 00 $0.0003$ $0.0006$
Temperature sensitivity
Table C.4 represents the stability of $\operatorname{atanh}$ and the SGA+ methods, expressed in true R-D loss, for different $\tau_{max}$ settings for the temperature schedule. As can be seen, the most optimal setting is with $\tau_{max}=1$ for each of the SGA+ methods. $\operatorname{atanh}$ obtains equal loss for $\tau_{max}\in[0.4,0.5]$. In general, we find that the linear method of SGA+ is least sensitive to changes in $\tau_{max}$ and has equal loss between $\tau_{max}\in[0.7,1]$. To further examine the stability of the linear function compared to $\operatorname{atanh}$, we subtract the best $\tau_{max}$, column-wise, from the linear and $\operatorname{atanh}$ of that column. We now see that the linear function is not only least sensitive to changes in $\tau_{max}$, but overall varies little compared to the best $\tau_{max}$ settings of the other methods. While the SSL has the largest drop in performance when reducing $\tau_{max}$, it achieves the highest performance overall for higher values of $\tau_{max}$.
@@ C.2 Analysis
Table C.1 presents the true R-D loss results for the interpolation with different $a$ settings for SSL for the mean-scale hyperprior model. In Figure C.7, the corresponding overall performance of the methods can be found. As can be seen in Figure 7(a), for $a=\{0.01,0.30\}$, the functions diverge, resulting in large loss values. For $a=0.65$, we find that the loss curve is slightly unstable at the beginning of training, which can be seen in the bending of the curve, indicating non-optimal settings. This may be due to the fact that we run all methods with the same $\tau_{max}=1$ for a fair comparison. Additionally, note that SSL with $a=0.65$ obtains a true R-D loss of $0.7410$ compared to $0.7418$ for $\operatorname{atanh}$ with the same settings. This is due to the fact that SSL, especially in the tails of the probability, is slightly more straight-curved compared to the $\operatorname{atanh}$ when looking at its probability space.
Table C.5: True R-D loss of two versus three-class rounding for SGA+ with the extended version of the linear, cosine, and SSL method at iteration 500 and in brackets after 2000 iterations.
@@
Three-class rounding
Table C.5 shows the true R-D loss for two- versus three-class rounding, at iteration $t=500$ and in brackets $t=2000$ iterations. For each method, we performed a grid search over the hyperparameters $r$ and $n$. Additionally, for the extended SSL, we also performed a grid search over $a$ and found the best setting to be $a=1.4$. As can be seen in the table, most impact is made with the extended version of the linear of SGA+, in terms of the difference between the two versus three-class rounding at iteration $t=500$ with loss difference $0.0035$ and $t=2000$ with $0.0006$ difference. There is a small difference at $t=500$ for the extended cosine version. In general, we find that running models longer results in convergence to similar values. SSL converges to equal values for two- and three-class rounding.
Semi-multi-rate behavior
Similar as in Appendix B.4, we experimented with different values for $\lambda$ to obtain a semi-multi-rate curve. For every pre-trained model, we ran SSL and $\operatorname{atanh}$ using $\lambda\in\{0.001,0.0025,0.005,0.01,0.02,0.04,0.08\}$. In Figure C.5, we have plotted the R-D curve of the base model (lime green line) and its corresponding R-D curves, obtained when refining the latents with the proposed $\lambda$’s for $\operatorname{atanh}$ and SSL. As can be seen, running the methods for $t=500$ iterations, SSL obtains best performance. While the longer you train, the closer together the performance will be.
@@
Figure C.7: Interpolation performance plots of different $a$ settings for SLL (a) True R-D Loss (b) Difference in loss (c) PSNR (d) BPP under the mean-scale hyperprior model.